I could use some help with this problem. I have tried solving it using the identities, but feel I am way off. Thank you in advance for your guidance!
cos(x/2) = -(sqrt3)/2

Answer by lwsshak3(11628)

cos(x/2) = -(sqrt3)/2
cos(x/2)=-√3/2
x/2=120˚
x=60˚